# Hao Chen1, Bo Zhang1, Chunhua Shen1 1

### v1

@@ 1 Introduction
Driven by the advent of generative modeling techniques [29, 24], there has been significant progress in video generation in the research community [6, 11, 2]. Tremendous research efforts have focused on adapting a pre-trained text-to-image diffusion model [6, 11] to a video generation model. Recently revolutionary leap has been made with the Sora [3] model which substantially scales a spatial-temporal transformer model. This model demonstrates remarkable video generation quality, dramatically extending the duration of generated videos from seconds to an unprecedented full minute. This technology profoundly reshapes the boundaries of generative AI, fostering anticipation towards hours-long movie production.
While predominant video generation methods adopt diffusion models, this paradigm is more suited for visual rendering and is less adept at modeling complex abstract logic and reasoning compared to autoregressive models, as evidenced in natural language processing. Moreover, diffusion models lack the flexibility to support arbitrary length. On the other hand, autoregressive models [38, 13, 12, 22] have shown superior ability in handling complex reasoning and are better at predicting what might happen in the next—a key element of a world model. Moreover, autoregressive models offer greater flexibility in handling varied lengths of predictions and can also benefit from mature training and inference infrastructure. As such, there are attempts that use the autoregressive model for video generation [14, 12]. However, autoregressive modeling is not as compute efficient as diffusion model for visual rendering and demands substantially more compute even for image rendering.
@@
Figure 1: The Titanic movie scenes generated by the proposed MovieDreamer. The generation is conditioned on the elaborate multimodal movie script. The original movie is unseen during training.
@@
Recent progress [8, 4, 25, 3] suggests that video generation quality in short duration (from seconds to a minute) can be consistently improved by scaling the compute due to the scaling law. In this work, we sidestep the short video generation problem and instead seek a generative framework for long video generation featuring complex narrative structures and intricate plot progressions, which is hardly achieved by merely scaling the current approaches. Our aim is to to push the boundaries of video generation beyond the limitations of short-duration content, enabling the creation of videos with rich, engaging storylines.
@@
In this paper, we introduce a hierarchical approach, dubbed MovieDreamer, that marries autoregressive modeling with diffusion-based rendering to achieve a novel synthesis of long-term coherence and short-term fidelity in visual storytelling. Our method leverages the strengths of autoregressive models to ensure global consistency in key movie elements like character identities, props, and cinematic styles, even amidst complex scene transitions. The model predicts visual tokens for specific local timespans, which are then decoded into keyframes and dynamically rendered into video sequences using diffusion rendering. Our method mirrors the traditional movie directing process, where intricate plots are decomposed into distinct, manageable scenes. Our method ensures that both the overarching narrative and the minute spatial-temporal details are maintained with high fidelity, thus enabling the automated production of movies that are both visually cohesive and contextually rich.
@@
To this end, we represent the whole movie using sparse keyframes and employ a diffusion autoencoder to tokenize each keyframe into compact visual tokens. We train an autoregressive model to predict the sequence of visual tokens. This model is initialized from a pretrained large language mode, a 7B LLaMA model [36], to leverage its rich world knowledge for better generalizability. As such, the model can be regarded as a multimodal model, which, rather than merely outputting the text, predicts the visual tokens with movie script conditioning.
@@
Three key designs need to be specially considered during the autoregressive training. One is to resort to various model and input perturbation techniques to combat the overfitting tendency since the high-quality movie training data is limited. Besides, we propose to condition the model using a novel multimodal script, which is structured hierarchically, including the rich description of the image style and scene elements, along with detailed text description of characters as well as their retrieved face embeddings. Such multimodal script facilitates narrative continuity across different segments of the movie while offering flexibility for character control. Moreover, we stochastically append a few reference frames ahead of the visual tokens of the same episode, and in this way the model acquires in-context learning capability and can synthesize high-quality personalized results. Such feature is particularly useful in generating coherent continuations of given movie episodes.
@@
We subsequently decode the predicted vision tokens using an image diffusion decoder and further render the image to the clip video using an image-to-video diffusion model. Importantly, we enhance this process by finetuning an identity-preserving diffusion decoder. This refinement effectively mitigates errors in vision token prediction and improves identity preservation in the resultant video clips. Our work is essentially orthogonal to the current efforts aimed at improving the short video generation quality via compute scaling and our method can benefit from these progresses. Our approach is validated through extensive experiments on a wide range of movie genres, showcasing an excelling ability to generate visually stunning and coherent long-form videos over the state-of-the-art.
@@
  • We introduce MovieDreamer, a novel hierarchical framework that marries autoregressive models with diffusion rendering to balance long-term narrative coherence with short-term visual fidelity. The method substantially extends the duration of generated video content to thousands of keyframes.
  • We generate the visual token sequence using a multimodal autogressive model. The autoregressive model supports zero-shot and few-shot personalized generation scenarios and supports variable length of keyframe prediction.
  • We use a novel multimodal script to hierarchically structures rich descriptions of scenes as well as the character’s identity. This approach not only facilitates narrative continuity across different segments of a video but also enhances character control and identity preservation.
  • Our method demonstrates superior generation quality with detailed visual continuity, high-fidelity visual details and the character’s identity-preserving ability.
@@ 2 Related Work
Video generation has advanced mainly through diffusion and autoregression methods. The popularization of Stable Diffusion has sparked extensive studies, improving video quality with various techniques [6, 11, 10, 7]. For long videos, strategies like [43, 25, 39, 4, 8] have proven effective. Despite their high computational load and challenges in managing complex sequences, diffusion models are widely used. Autoregression, on the other hand, is less common but excels in specific areas like autonomous driving by employing unified token spaces [12], although it struggles with continuity in complex scenes. Our approach combines the strengths of both methods in a hierarchical and scalable manner, enhancing control and flexibility irrespective of video length.
@@
Recent advancements lead to the emergence of visual language models (VLMs) through enabling large language model to comprehend visual modalities [20, 36].VLMs have demonstrated their power in generating images [22, 34, 33], in addition to language. Moreover, ongoing research [46, 18, 16, 47] has showcased the robust capabilities of Multimodal Language Models (MLLMs) in comprehending video content. These advancements have solidified the effectiveness of VLMs in addressing complex problems. Building on this foundation, our approach leverages VLMs to facilitate the generation of extremely long video content.
@@
The crux of visual story generation lies in ensuring consistency across generated images. Similar to video generation, diffusion and autoregression are also two primary manners. For diffusion-based methods, existing methods predominantly employ conditional generation to maintain this coherence  [31, 40, 35, 9]. Some [17, 41] have shown promising results in retaining high-quality facial features and identity. Other approaches are based on the autoregressive ideas  [5, 19, 27, 23]. However, these approaches generally yield mediocre results and are challenging to scale up.Story and video generation share intrinsic links, with overlapping solutions. Our method uniquely integrates these, offering a unified approach.
@@ 3.1 Overview
We propose a novel framework for generating extended video sequences that leverages the strengths of autoregressive models for long-term temporal consistency and diffusion models for high-quality image rendering. Our method takes multimodal scripts as conditions, predicts keyframe tokens in an autoregressive manner, and uses these frames as anchors to produce full-length videos. Our method offers flexibility to supports zero-shot generation along with few-shot scenarios in which the generation results are required to follow the given style. We take special care of preserving the identity of the chracters throughout the multimodal script design, autoregressive training and diffusion rendering. We illustrate the overall framework in Figure 2.
@@ 3.2 Keyframe Tokenization via Diffusion Autoencoder
To create a concise yet faithful representation of images, we employ a diffusion autoencoder. Our encoder, $\mathcal{E}$, contains a pretrained CLIP vision model [26] and a transformer-based token compressor, which encodes an image $\mathbf{x}$ into reduced number of compressed tokens denoted by $\mathbf{e}$. The decoder, $\mathcal{D}$, uses these tokens to reconstruct the image, which adopts pretrained SDXL diffusion model and yields an 896$\times$512 reconstructed image. The extracted latent image tokens are fed into the diffusion decoder via cross-attention. We train the compressor and the decoder while leaving the CLIP vision model frozen. The loss for training this diffusion autoencoder is as follows:
@@
where $\mathbf{x}_{0}$ is the input image and $\mathbf{z}_{t}$ is its corresponding noisy latent at timestep $t$, i.e.$\mathbf{z}_{t}=\alpha_{t}\mathbf{x}+\sigma_{t}\bm{\epsilon}$, where $\bm{\epsilon}\in\mathcal{N}(0,I)$, and $\alpha_{t}$ and $\beta_{t}$ define a noising schedule. In our experiment, we empirically find that merely two tokens can sufficiently characterize major semantics of keyframes, corroborating the finding of previous work [42].
@@ 3.3 Autoregressive Keyframe Token Generation
xWe initialize our autoregressive model with a pretrained large language model, LLaMA2-7B [36]. In contrast to LLMs, our model $\mathcal{G}$ predicts the compressed vision tokens based on the input multi-modal scripts $\mathbf{M}$ and history information $\mathbf{H}$, which is formulated as $\mathbf{e}_{t}=\mathcal{G}(\mathbf{H}_{<t},\mathbf{M}_{t})$.
We initiate our approach by leveraging a pretrained large language model, specifically LLaMA2-7B [36], to construct our autoregressive model, $\mathcal{G}$. Unlike traditional LLMs, $\mathcal{G}$ is designed to predict compressed vision tokens from multimodal scripts $\mathbf{M}$ and historical data $\mathbf{H}$. The model function is defined as $\mathbf{e}_{t}=\mathcal{G}(\mathbf{H}_{<t},\mathbf{M}_{t})$.
Traditional LLMs often employ cross-entropy loss for training, which is suitable for discrete outputs. However, our model deals with continuous real-valued image tokens, rendering cross-entropy inapplicable.
Inspired by the GIVT [37], we adopt a $k$-mixture Gaussian Mixture Model (GMM) to effectively model the distribution of these real-valued tokens. This involves parameterizing the GMM with $2kd$ means, $2kd$ variances, and $k$ mixing coefficients.
These parameters are obtained from a modified linear output layer of the autoregressive model, enabling sampling of continuous tokens from the GMM. The model is trained by minimizing the negative log-likelihood:
@@
It is challenging to massively acquire large amount of high-quality, long-duration video data. We manage to collect 50k long videos, yielding 5M keyframes for autoregressive training. Yet this data volume is relatively limited, leading to a severe overfitting issue where the generative model fails to generalize well during testing. To address this, we implement several strategies which are crucial for training:
@@
We develop a well-structured multi-modal script format to serve as input for autoregressive models, as depicted in Figure 14. Our script integrates multiple dimensions: characters, scene elements, and narrative arcs. Accurately representing character appearance using text alone is challenging; therefore, we combine textual descriptions with facial embeddings to provide a more detailed representation of each character. To facilitate the processing by the autoregressive model, we structure the script format to distinctly delineate these elements.
For non-textual modalities such as facial embeddings and compressed tokens are projected to LLaMA’s embedding space using multi-layer perceptron. The primary challenge arises with textual data, which tends to produce long sequences, thereby consuming excessive token space and limiting the model’s contextual breadth. To mitigate this, we treat text as a separate modality, dividing it into “identifiers” and “descriptions” (see Figure 2). Identifiers are concise statements that establish the script’s structure. In contrast, descriptions, which detail the attributes for generation, are each encoded into a single [CLS] token using CLIP and then projected into the unified input space.
This method significantly extends the usable contextual length during training by condensing entire sentences into single tokens. We utilize LongCLIP [44] as our text encoder for descriptions, supporting inputs up to 248 tokens, which enhances our capacity to handle detailed narrative content. Consequently, the multi-modal script at timestep $t$ and preceding historical data are represented as:
@@
where $\mathbf{d}_{t}$, $\mathbf{i}_{t}$, and $\mathbf{c}_{t}$ denote the embeddings for descriptions, identifiers, and characters’ facial embeddings, respectively. $\mathbf{e}_{<t}$ represents the previously predicted compressed frame tokens. The negative log-likelihood loss is then formulated as:
@@
Our model is versatile, supporting both zero-shot and few-shot generation modes. In zero-shot mode, the model generates content from a text prompt alone. In few-shot mode, it leverages a small set of user-provided reference images to align the generated content more closely with user preferences, without necessitating further training. This capability ensures that users can efficiently produce high-quality, customized visual content that aligns with their desired themes and styles.
3.4 ID-preserving Diffusion Rendering
In our approach, while the primary diffusion decoder $\mathcal{D}$ adeptly reconstructs target images, it occasionally falls short in capturing fine-grained details, notably in facial features, due to the attenuation of details in compressed tokens. To address this, we enhance the cross-attention module within $\mathcal{D}$. This enhancement involves the integration of both descriptive text embeddings, $\mathbf{d}$, and face embeddings, $\mathbf{f}$, derived from the multimodal script.

### v2

# MovieDreamer: Hierarchical Generation for Coherent Long Visual Sequences

@@ 1 Introduction
Driven by the advent of generative modeling techniques (Rombach et al., 2022; Podell et al., 2023), there has been significant progress in video generation in the research community (Guo et al., 2024; Ho et al., 2022b; Blattmann et al., 2023). Tremendous research efforts have focused on adapting a pre-trained text-to-image diffusion model (Guo et al., 2024; Ho et al., 2022b) to a video generation model. Recently revolutionary leap has been made with the Sora (Brooks et al., 2024) model which substantially scales a spatial-temporal transformer model. This model demonstrates remarkable video generation quality, dramatically extending the duration of generated videos from seconds to an unprecedented full minute. This technology profoundly reshapes the boundaries of generative AI, fostering anticipation towards hours-long movie production.
While predominant video generation methods adopt diffusion models, this paradigm is more suited for visual rendering and is less adept at modeling complex abstract logic and reasoning compared to autoregressive models, as evidenced in natural language processing. Moreover, diffusion models lack the flexibility to support arbitrary length. On the other hand, autoregressive models (Wang et al., 2023a; Jiang et al., 2022; Hu et al., 2023; Pan et al., 2023) have shown superior ability in handling complex reasoning and are better at predicting what might happen in the next—a key element of a world model. Moreover, autoregressive models offer greater flexibility in handling varied lengths of predictions and can also benefit from mature training and inference infrastructure. As such, there are attempts that use the autoregressive model for video generation (Kondratyuk et al., 2023; Hu et al., 2023). However, autoregressive modeling is not as compute efficient as diffusion model for visual rendering and demands substantially more computing even for image rendering.
@@
Figure 1: The Titanic movie scenes generated by the proposed MovieDreamer. The generation is conditioned on the elaborate multimodal movie script. The original movie is unseen during training. Please refer to Figure 1011121314 for more results.
@@
Recent progress (Henschel et al., 2024; Chen et al., 2023; Qiu et al., 2023; Brooks et al., 2024) suggests that video generation quality in short duration (from seconds to a minute) can be consistently improved by scaling the compute due to the scaling law. In this work, we sidestep the short video generation problem and instead seek a generative framework for long video generation featuring complex narrative structures and intricate plot progressions, which is hardly achieved by merely scaling the current approaches. Our aim is to push the boundaries of video generation beyond the limitations of short-duration content, enabling the creation of videos with rich, engaging storylines.
@@
In this paper, we introduce a hierarchical approach, dubbed MovieDreamer, that marries autoregressive modeling with diffusion-based rendering to achieve a novel synthesis of long-term coherence and short-term fidelity in visual storytelling. Our method leverages the strengths of autoregressive models to ensure global consistency in key movie elements like character identities and cinematic styles, even when the camera switches back and forth. The model predicts visual tokens for specific local timespans, which are then decoded into keyframes and dynamically rendered into video sequences using diffusion rendering. Our method mirrors the traditional movie directing process, where intricate plots are decomposed into distinct, manageable scenes. Our method ensures that both the overarching narrative and the minute spatial-temporal details are maintained with high fidelity, thus enabling the automated production of movies that are both visually cohesive and contextually rich.
@@
To this end, we represent the whole movie using sparse keyframes and employ a diffusion autoencoder to tokenize each keyframe into compact visual tokens. We train an autoregressive model to predict the sequence of visual tokens. This model is initialized from a pretrained large language mode, a 7B LLaMA model (Touvron et al., 2023), to leverage its rich world knowledge for better generalizability. As such, the model can be regarded as a multimodal model, which, rather than outputting the text, predicts the visual tokens with movie script conditioning.
@@
Three key designs need to be specially considered during the autoregressive training. One is to resort to various model and input perturbation techniques to combat the overfitting tendency since the high-quality movie training data is limited. Besides, we propose to condition the model using a novel multimodal script, which, for each keyframe, includes the rich description of the image style and scene elements, along with detailed text description of characters as well as their retrieved face embeddings. Such multimodal script facilitates narrative continuity across different segments of the movie while offering flexibility for character control. Moreover, we stochastically append a few reference frames of the same episode at the beginning of the input, and in this way, the model acquires in-context learning capability and can synthesize higher-quality personalized results. Such feature is particularly useful in generating coherent continuations of given movie episodes.
@@
We subsequently decode the predicted vision tokens using an image diffusion decoder and further render the image to the video clip using an image-to-video diffusion model. Importantly, we enhance this process by finetuning an identity-preserving diffusion decoder. This refinement effectively improves identity preservation in the resultant video clips. Our work is essentially orthogonal to the current efforts aimed at improving the short video generation quality via compute scaling and our method can benefit from these progresses. Our approach is validated through extensive experiments on a wide range of movie genres, showcasing an excelling ability to generate visually stunning and coherent long-form videos over the state-of-the-art.
@@
  • We introduce MovieDreamer, a novel hierarchical framework that marries autoregressive models with diffusion rendering to generate long-term coherent stories and videos with multiple characters well-preserved. The method substantially extends the duration of generated video content to thousands of keyframes. To the best of our knowledge, we are the first to tackle the ultra-long narrative video generation problem, rather than merely clip looping.
  • We generate the visual token sequence using a multimodal autoregressive model. The autoregressive model supports zero-shot and few-shot personalized generation scenarios and supports variable length of keyframe prediction.
  • We use a novel multimodal script to hierarchically structure rich descriptions of scenes, plots, as well as the character’s identity. This approach not only facilitates narrative continuity across different segments of a video but also enhances character control and identity preservation.
  • Our method demonstrates superior generation quality with detailed visual continuity, high-fidelity visual details, and the character’s identity-preserving ability. Both story generation and video generation can be addressed with our method.
@@ 2 Related Work
Video generation has advanced mainly through diffusion and autoregression methods. The popularization of Stable Diffusion (Rombach et al., 2022) has sparked extensive studies, improving video quality with various techniques (Guo et al., 2024; Ho et al., 2022b; a; He et al., 2023). For long videos, strategies like (Yin et al., 2023; Qiu et al., 2023; Wang et al., 2023b; Chen et al., 2023; Henschel et al., 2024) can only generate simple results. Despite their high computational load and challenges in managing complex sequences, diffusion models are widely used. Autoregression, on the other hand, is less common but excels in specific areas like autonomous driving by employing unified token spaces (Hu et al., 2023), although it struggles with continuity in complex scenes. Our approach combines the strengths of both methods in a hierarchical and scalable manner, enhancing control and flexibility irrespective of video length.
@@
Recent advancements in NLP lead to the emergence of visual language models (VLMs) through enabling large language model (Touvron et al., 2023) to comprehend visual modalities (Liu et al., 2023b; Bai et al., 2023a; b; Chen et al., 2024). In addition to language, VLMs have demonstrated their power in generating images (Pan et al., 2023; Sun et al., 2023c; b). Moreover, ongoing research (Zhang et al., 2023; Lin et al., 2023; Li et al., 2023a; Zhao et al., 2022) has showcased the robust capabilities of Multimodal Language Models (MLLMs) in comprehending video content. Some works (Cai et al., 2024; Cheng et al., 2024; Ren et al., 2024b; Rasheed et al., 2024; Lai et al., 2024) enable VLMs to assist in spatial understanding and other visual tasks. VLMs (Jiang et al., 2022; Kim et al., 2024; Etukuru et al., 2024; Brohan et al., 2023; Driess et al., 2023) also demonstrate the powerful capabilities in robotics. These advancements have solidified the effectiveness of VLMs in addressing complex problems. Building on this foundation, our approach leverages VLMs to facilitate the generation of extremely long video content.
@@
The crux of visual story generation lies in ensuring consistency across generated images. Similar to video generation, diffusion and autoregression are also two primary manners. Existing diffusion-based methods predominantly employ conditional generation to maintain the coherence  (Su et al., 2023; Wang et al., 2023c; Tewel et al., 2024; Hertz et al., 2023). Some (Li et al., 2023b; Wu et al., 2024) have shown promising results in retaining high-quality facial features and identity. Other approaches are based on the autoregressive ideas  (Feng et al., 2023; Liu et al., 2023a; Rahman et al., 2023; Pan et al., 2022). However, these approaches generally yield mediocre results and are challenging to scale up. Story and video generation share intrinsic links, with overlapping solutions. Our method uniquely integrates these, offering a unified approach.
@@ 3.1 Overview
We propose a novel framework for generating extended video sequences that leverages the strengths of autoregressive models for long-term temporal consistency and diffusion models for high-quality image and video rendering. Our method takes multimodal scripts as conditions, predicts keyframe tokens in an autoregressive manner, renders the tokens into images and uses these images as anchors to produce full-length videos. Our method offers flexibility to support zero-shot generation along with few-shot scenarios in which the generation results are required to follow the given reference. We take special care of preserving the identity of the characters throughout the multimodal script design, autoregressive training, and diffusion rendering. We illustrate the overall framework in Figure 2.
@@ 3.2 Keyframe Tokenization via Diffusion Autoencoder
To create a concise yet faithful representation of images, we employ a diffusion autoencoder. Our encoder, $\mathcal{E}$, contains a pretrained CLIP vision model (Radford et al., 2021) and a transformer-based token compressor, which encodes an image $\mathbf{x}$ into a reduced number of compressed tokens denoted by $\mathbf{e}$. The decoder, $\mathcal{D}$, is finetuned based on the pretrained SDXL and yields an 896$\times$512 reconstructed image by leveraging the compressed tokens in the cross-attention module. We train the compressor and the decoder while leaving the CLIP vision model frozen. The loss for training this diffusion autoencoder is as follows:
@@
where $\mathbf{x}_{0}$ is the input image and $\mathbf{z}_{t}=\alpha_{t}\mathbf{x}+\sigma_{t}\bm{\epsilon}$ is its corresponding noisy latent at timestep $t$. $\bm{\epsilon}\in\mathcal{N}(0,I)$ is the Gaussian noise. $\alpha_{t}$ and $\sigma_{t}$ define the noise schedule. In our experiment, we empirically find that merely two tokens can sufficiently characterize major semantics of keyframes, corroborating the finding of previous work (Yang et al., 2022).
3.3 ID-preserving Diffusion Rendering
In our approach, while the primary diffusion decoder $\mathcal{D}$ adeptly reconstructs target images, it occasionally falls short in capturing fine-grained details, notably in facial features, due to the attenuation of details in compressed tokens. To address this, we enhance the cross-attention module within $\mathcal{D}$. This enhancement involves the integration of both descriptive text embedding, $\mathbf{d}$, and face embedding, $\mathbf{f}$, derived from the multimodal script. Specifically, the face embedding and description embedding is projected to the same dimension as the compressed token with two additional MLPs, which are then concatenated together and serve as the input to the keys and values of cross-attention modules in $\mathcal{D}$.
@@ 3.4 Autoregressive Keyframe Token Generation
We initiate our approach by leveraging a pretrained large language model, specifically LLaMA2-7B (Touvron et al., 2023), to construct our autoregressive model, $\mathcal{G}$. Unlike traditional LLMs, $\mathcal{G}$ is designed to predict compressed vision tokens from the multimodal script $\mathbf{M}$ of the corresponding keyframe and historical information $\mathbf{H}$, which is formulated as $\mathbf{e}_{t}=\mathcal{G}(\mathbf{H}_{<t},\mathbf{M}_{t})$.
Traditional LLMs often employ cross-entropy loss for training, which is suitable for discrete outputs. However, our model deals with continuous real-valued image tokens, which makes cross-entropy inapplicable. Inspired by  Tschannen et al. (2024), we adopt a $k$-mixture Gaussian Mixture Model (GMM) to effectively model the distribution of these real-valued tokens. This involves parameterizing the GMM with $2kd$ means, $2kd$ variances, and $k$ mixing coefficients. These parameters are predicted by attaching an additional trainable linear layer at the end of the autoregressive model, enabling the construction of the GMM and the sampling of continuous tokens from the GMM. The model is trained by minimizing the negative log-likelihood:
@@
It is challenging to massively acquire large amounts of high-quality, long-duration video data. We managed to collect 50k long videos, yielding 5M keyframes for autoregressive training. Yet this data volume is relatively limited, leading to a severe overfitting issue where the generative model fails to generalize well during testing. To address this, we implement several strategies which are crucial for training:
@@
We develop a well-structured multi-modal script format to serve as input for autoregressive models, as depicted in Figure 16. Our script integrates multiple dimensions: characters, scene elements, and plots. Accurately representing character appearance using text alone is challenging; therefore, we combine textual descriptions with facial embeddings (Zheng et al., 2022) to provide a more detailed representation of each character. To facilitate the processing by the autoregressive model, we structure the script format to distinctly delineate these elements.
For non-textual modalities such as facial embeddings and compressed tokens are projected to LLaMA’s embedding space using multi-layer perceptron. The primary challenge arises with textual data, which tends to produce long sequences, thereby consuming excessive token space and limiting the model’s contextual breadth. This challenge poses a significant obstacle to achieving our goal of long content generation. To mitigate this, we treat text as a separate modality, dividing it into “identifiers” and “descriptions” (elaborated in Appendix A). Identifiers establish the script’s structure, while descriptions detail the attributes for generation. Descriptions are segmented into sentences, with each encoded into a single [CLS] token using CLIP and then projected into the unified input space. In other words, we perform the sentence-level tokenization.
This method significantly extends the usable contextual length during training by condensing an entire sentence into a single token. We utilize LongCLIP (Zhang et al., 2024) as our text encoder for descriptions, supporting inputs up to 248 tokens, which enhances our capacity to handle detailed narrative content. Consequently, the multi-modal script at timestep $t$ and preceding historical information are represented as:
@@
where $\mathbf{d}_{t}$, $\mathbf{i}_{t}$, and $\mathbf{c}_{t}$ denote the embeddings for descriptions, identifiers, and characters’ facial embeddings, respectively. $\mathbf{e}_{<t}$ represents the previously predicted compressed frame tokens. The negative log-likelihood loss in Equation 2 is then formulated as:
@@
Our model is versatile, supporting both zero-shot and few-shot generation modes. In zero-shot mode, the model generates content from the multimodal script. In few-shot mode, it leverages a small set of user-provided reference images to align the generated content more closely with user preferences, without necessitating further training. This capability ensures that users can efficiently produce high-quality, customized visual content that aligns better with their desired target.